OpenStudy (anonymous):

The area of the base of a prism is 50 mm2. The perimeter of the base is 30 mm. The height of the prism is 7 mm. What is the surface area of the prism? ____________________mm(2)

OpenStudy (anonymous):

I am not sure, I will give medal and fan to whom ever knows and explains their answer

OpenStudy (anonymous):

@bibby @sleepyjess @Somy @Sparklestaraa @Starbucks @Jaynator495 @love10129151

jimthompson5910 (jim_thompson5910):

lateral surface area = (perimeter of base)*(height of prism)

OpenStudy (anonymous):

plz help

jimthompson5910 (jim_thompson5910):

total surface area = 2*(area of 1 base) + (lateral surface area)

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

30x7

OpenStudy (anonymous):

30x7=210

jimthompson5910 (jim_thompson5910):

210 is your lateral surface area

OpenStudy (anonymous):

50x210=10500

OpenStudy (anonymous):

10500 is the answer??

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (anonymous):

really, that was easier than i thought

OpenStudy (anonymous):

okay, then what is i am confused

jimthompson5910 (jim_thompson5910):

total surface area = 2*(area of 1 base) + (lateral surface area) total surface area = 2*(50) + (210) total surface area = ??

OpenStudy (anonymous):

OOOOOHHHH

OpenStudy (anonymous):

we are not multiplieing, we are adding

OpenStudy (anonymous):

then the answer is 260, right?!

OpenStudy (anonymous):

right @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no I don't know how you're getting 260

jimthompson5910 (jim_thompson5910):

total surface area = 2*(area of 1 base) + (lateral surface area) total surface area = 2*(50) + (210) total surface area = ??

jimthompson5910 (jim_thompson5910):

I think you forgot about the 2 out front

OpenStudy (anonymous):

310

jimthompson5910 (jim_thompson5910):

yep 310

OpenStudy (anonymous):

thaks, can u help me with some more

OpenStudy (anonymous):

i will tag u

jimthompson5910 (jim_thompson5910):

sure I'll do 1 more